# Investigating Efficacy of Perplexity in Detecting LLM-Generated Code

### v1

@@ VI-B For Tool Development and Usages
When detecting LLMgCode from HaCode, we should first detect whether it is LLMgCode or whether it is entirely or partially generated by LLMs, then detect which parts (lines or blocks) are generated by LLMs.
@@
The first step can be formulated as a classification problem. However, the experimental results show that none of the three detection methods (classifiers) are effective. Perplexity only performs well for detecting code in C/C++ (from the perspective of language), and large (from the perspective of scale). The feature-based and pre-training-based methods have high effectiveness but poor generalization capability. Overall, they are all acceptable with an AUC of more than 70%. An ensemble learning approach can make full use of each separate method’s advantages meanwhile overcoming disadvantages to some extent. Taking detecting C code as an example, we recommend using DetectCodeGPT (Perplexity), Bukhari (feature-based method), and GPTSniffer (pre-training-based) to develop an ensemble detection tool. By employing an ensemble learning strategy such as MoE (Mixture Of Experts), the detection tool can automatically learn which method(s) to prioritize under different contexts.
@@
The second step is fine-grained detection, i.e., identifying which lines are generated by LLMs. Having understood the code is generated by LLMs, the second step is to first calculate the perplexity of each token in the code and then calculate the average perplexity for each line. These lines with low perplexity are detected as LLMgCode. We recommend highlighting LLMgCode and HaCode in different colors to remind developers.
The detection tool can be applied in but not limited to the following scenarios.
  • Software Supply Chain Management: LLMgCode has become a new third-party asset in software supply chain development [15]. The detection tool helps enterprises manage LLMgCode for tracking accountabilities, checking violations, etc.
  • Code Review: Since LLMgCode is more prone to include defects and vulnerabilities, attention should be placed on it during code review. The detection tool is useful for highlighting LLMgCode to remind reviewers to perform additional checks such as plagiarism detection.
  • Productivity Assessment: Many developers use LLMs to assist programming, making traditional LOC-based productivity assessment less useful. The detection tool helps enterprises assess developers’ productivity more accurately than before.
  • Originality Assessment: LLMgCode is sometimes not considered original. The detection tool is useful for situations where originality is emphasized, including but not limited to software competitions and student assignments.
VII Limitation
@@
This section discusses the possible limitations of this work and our efforts to mitigate their effects.
Test Scale
Due to the budget and time costs, we randomly tested 500 samples for each group in each experiment. For instance, the groups from the perspective of difficulty are easy, middle, and hard. The scale of our testing for each group is comparable to previous related work [14, 16]. Our testing scale is considered acceptable to mitigate biases introduced by random sampling.
Subject Selection
For the experimental subjects, we included several detection methods that were not originally proposed to detect LLMgCode, e.g., DetectGPT and NPR. However, they were proposed to detect LLM-generated text. As source code can be considered as one specific type of text, we believe these detection methods are also suitable for detecting LLMgCode. This setting was also used in previous related work [14, 16]. Moreover, although Perplexity is an unsupervised method, we compared it against supervised methods because Shi et al. [14] reported that Perplexity outperformed supervised methods. We excluded methods whose implementation details are unclear. For example, we excluded AIgCode Detector [26] because it did not specify how perturbations are performed.
Dataset Diversity
For the diversity of HaCode, we collected HaCode from CodeNet only. CodeNet collects HaCode solutions from 14K submits. The HaCode solutions in CodeNet are in 55 languages. 54% of HaCode solutions in CodeNet are accepted, 30% are marked as wrong solutions, and the remaining suffer from one of the possible rejection causes. The richness of CodeNet guarantees the diversity of HaCode in our experiments. For the diversity of LLMgCode, we only evaluated Perplexity in detecting LLMgCode generated by GPT-3.5 in $\text{RQ}_{1}$, which may bias experimental results. However, in $\text{RQ}_{2}$ and $\text{RQ}_{3}$, we found that Perplexity achieved similar results in detecting LLMgCode generated by GPT-4o, Gemini, and Llama. The results show that using code generated by GPT-3.5 is acceptable.
VIII Conclusion
@@
LLMgCode has become increasingly common in software development; however, its security and compliance remain challenging. Moreover, LLMgCode is often mixed with HaCode. Human-LLM code authorship attribution is not only important to improve the transparency of software development, but also to increase the trust of third-party software in the supply chain. In this article, we conduct a large-scale investigation of Perplexity, the state-of-the-art method for detecting LLMgCode from HaCode. Toward realistic evaluations, we use 11,664 HaCode snippets and 13,164 LLMgCode snippets to evaluate Perplexity from three perspectives: detection accuracy, generalization capability, and inference efficiency. The experimental results show that Perplexity is less effective than other methods. Perplexity only performs well in C/C++ from the perspective of language and in large-scale from the perspective of scale. Additionally, Perplexity has a slower inference speed than other methods. However, Perplexity is still the most promising method because it has the best generalization capacity. Furthermore, Perplexity has good interpretability and can be applied for fine-grained detection. In the future, we are going to continually improve Perplexity to detect LLMgCode from HaCode and evaluate it in industry.

### v2

@@ 6.3. For Tool Development and Usages
When detecting LLMgCode, it is recommended to first detect whether it is LLMgCode and whether it is entirely or partially generated by LLMs, then detect which parts (lines or blocks) are generated by LLMs.
@@
The first step can be formulated as a classification problem. However, the experimental results show that none of the three detection methods (classifiers) are effective. Perplexity only performs well for detecting code in C/C++ (from the perspective of language), and with a large scale (from the perspective of scale). The feature- and pre-training-based methods have high effectiveness but poor generalization capability. Overall, they are all acceptable with an AUC of more than 70%. An ensemble learning approach can fully use each separate method’s advantages while overcoming its disadvantages to some extent. Taking detecting C code as an example, we recommend using DetectCodeGPT (Perplexity), Bukhari (feature-based method), and GPTSniffer (pre-training-based) to develop an ensemble detection tool. By employing an ensemble learning strategy such as MoE (Mixture Of Experts), the detection tool can automatically learn which method(s) to prioritize in different contexts.
@@
The second step is fine-grained detection, i.e., identifying which lines are generated by LLMs. Having understood that the code is generated by LLMs, the second step is to first calculate the perplexity of each token in the code and then calculate the average perplexity for each line. These lines with low perplexity are detected as LLMgCode. We recommend highlighting LLMgCode and HaCode in different colors to remind developers.
The detection tool can be applied in but not limited to following scenarios.
Software supply chain management. LLMgCode has become a new third-party asset in software supply chain development (Bukhari et al., 2023). The widespread use of LLMgCode raises concerns regarding code provenance, accountability, and the potential for introducing bugs or vulnerabilities into software. To address these challenges, the detection tool provides a solution for managing LLMgCode. By detecting the proportion of LLMgCode in third-party libraries, enterprises can assess the reliability of these libraries. By detecting LLMgCode in internal code and centrally managing it as a third-party asset, enterprises can verify compliance with licensing requirements to reduce violations. When there is a problem with the code generated by a certain LLM, it can be quickly located and measures can be taken.
Code review. Since LLMgCode is more prone to include defects (Perry et al., 2024; Sandoval et al., 2023), particular attention is required during the code review process. LLMgCode can lead to unintended risks due to the inherent limitations of LLMs, such as the inclusion of insecure code in the training set (Asare et al., 2023). Therefore, it is crucial for reviewers to be aware of the potential risks associated with LLMgCode and to perform extra scrutiny. The detection tool serves as a valuable aid in this process by automatically detecting and highlighting LLMgCode, ensuring that reviewers are reminded to apply additional checks. These checks may include not only the detection of unique vulnerability patterns targeting LLMgCode, but also plagiarism detection to ensure that LLMgCode is compliant with licensing regulations.
Productivity assessment. Many developers use LLMs to assist with programming, which has led to a shift in how developer productivity is measured (Shi et al., 2025). Traditional metrics, such as LOC, have become less applicable because a significant portion of the submitted code is LLMgCode. LOC-based assessments treat a line of LLMgCode and HaCode as equivalent, resulting in biased evaluations of developers’ actual contributions. To mitigate this challenge, the detection tool helps detect LLMgCode in the submitted code, enabling enterprises to apply more detailed and approximate criteria for assessment. After distinguishing LLMgCode and HaCode, custom evaluations can be performed for each, allowing for a fairer assessment of developer productivity.
Originality assessment. LLMgCode is sometimes not considered original (Nguyen et al., 2024), which raises a concern in contexts that demand independent problem-solving, intellectual property protection, or academic integrity. The detection tool is useful in such situations by automatically determining whether the submitted code is mainly written by humans or generated by LLMs. It can be applied in various scenarios, including but not limited to software competitions, where independent problem-solving skills are valued, and student assignments, where academic honesty and plagiarism prevention are essential. Using this tool, organizations can ensure that submissions meet the necessary standards of originality.
7. Threat to Validity
@@
In this section, we discuss the possible threats to validity and our efforts to mitigate their effects.
Internal Validity: The functions in our dataset are designed to solve specific programming puzzles and are mainly self-contained. This may differ from functions in real-world software repositories, where many customized functions exist and cross-file function calls are common. Although we can extract HaCode from repositories (e.g., GitHub), generating accurate task descriptions from HaCode is difficult without domain knowledge. Without task descriptions, LLMs can not generate precise LLMgCode solutions. Using HaCode and LLMgCode which solve different problems can introduce bias. For example, using HaCode that addresses real-world tasks and LLMgCode that solves programming puzzles may result in LLMgCode being less complex than HaCode, causing bias in the evaluation results. Therefore, we build our dataset based on CodeNet (Puri et al., 2021), which offers a clear task description for each problem and has been widely used in prior work (Xu and Sheng, 2024; Naik et al., 2023; Pan et al., 2024).
Construct Validity: We excluded trivial problems in the dataset as they rarely occur in real-world scenarios. According to  Sheard et al. (2013), there is a strong correlation between code length and problem difficulty. Therefore, we excluded programming problems whose human solutions are all less than 30 LOC to filter out trivial problems. We adopted the thresholds used in prior work–10 (Xu and Sheng, 2024) and 100 (Nguyen et al., 2024)–to determine the initial range for our threshold setting. Within this range, we tested the number of filtered problems under different thresholds at intervals of 10. When the threshold exceeds 50, the number of filtered problems increases rapidly, and this growth becomes steep when the threshold exceeds 80. Excessive filtering introduces risks of excluding non-trivial problems. Therefore, to ensure the generalization of threshold setting, we adopt a safe value of 30. We manually checked the problems filtered by the threshold of 30. All of them are trivial problems or problems without valid submissions.
External Validity: The models used for perplexity calculation range from 1 billion to 7 billion parameters. We examine the statistical correlation between the model size and the evaluation results using Spearman’s rank correlation (Spearman, 1961), which has been widely used in previous work (Niazi, 2015; Hijazi et al., 2022; Lu et al., 2012; Zhou et al., 2009). The resulting p-values are all above 0.05, indicating no statistically significant correlation. This can be explained by the fact that, although the models for perplexity calculation differ in parameter size, they are all much smaller than the code-generating models (such as 175 billion). As a result, their code understanding capabilities are all significantly weaker than those of generation models, resulting in no statistically significant correlation between model size and evaluation results. Small models generally have weaker code understanding capabilities than large models. However, it is impractical for us to locally deploy large generation models (e.g., GPT-3.5) due to resource constraints. Therefore, we use the same model as prior work (Shi et al., 2025) when calculating perplexity.
8. Conclusion
@@
LLMgCode has become increasingly common in daily software development; however, its quality and security remain challenging. Moreover, LLMgCode is sometimes mixed with HaCode. Human-LLM code authorship attribution improves the transparency of software development and increases the trust of third-party software in the supply chain. In this article, we conduct a large-scale investigation of Perplexity, the state-of-the-art method for detecting LLMgCode from HaCode. Toward realistic evaluations, we use 11,664 HaCode snippets and 13,164 LLMgCode snippets to evaluate Perplexity from three perspectives: detection accuracy, detection speed, and generalization capability. The experimental results show that Perplexity is less effective than other methods. Perplexity only performs well in C/C++ from the perspective of language and in a large scale from the perspective of scale. Additionally, Perplexity has a slower detection speed than other methods. However, Perplexity demonstrates strong generalization capability, making it robust in scenarios where the LLM source is unknown. It also has good interpretability and can be naturally applied to fine-grained detection. In summary, there is a trade-off between detection accuracy, speed, and generalization capability: Perplexity achieves similar performance across LLMs but has relatively low accuracy and slow speed. While supervised methods achieve higher accuracy and faster speed, their performance varies when the LLMs for code generation are unknown. In the future, we will continually improve Perplexity to detect LLMgCode from HaCode and evaluate it in industry.
Data Availability
The data and materials of this article are available at https://figshare.com/s/10035be3b64e30e6e779.
Acknowledgments
This work is supported by the Key Research and Development Program of Jiangsu Province (No.BE2021002-2), the Natural Science Foundation of Jiangsu Province (No.BK20241195), the Open Project of Key Laboratory of Industry and Information Technology Ministry for Software Fusion Application and Testing Verification, the National Natural Science Foundation of China (No.62302210, No.62202219), and the Innovation Projects and Overseas Open Projects of State Key Laboratory for Novel Software Technology (Nanjing University) (ZZKT2025A12, ZZKT2025B18, ZZKT2025B20, ZZKT2025B22, KFKT2025A17, KFKT2025A19, KFKT2025A20, KFKT2024A02, KFKT2024A13, KFKT2024A14, KFKT2023A09, KFKT2023A10). Alberto Bacchelli gratefully acknowledges the support of the Swiss National Science Foundation through the SNF Project 200021_197227.